# references.bib

### v1

@@ 1 Introduction
Ramsey theory is an area of mathematics asserting that each sufficiently large object contains a still large well-structured subobject. The foundational theorem, proved by Frank Ramsey in 1930, states that every sufficiently large graph contains a still large homogeneous subgraph, i.e. a clique or an independent set [Ramsey1930]. It is well known by a theorem of Erdős and Szekeres [Erdos1935] and a probabilistic construction by Erdős [Erdos1947] that there is always a logarithmically large homogeneous set, but we cannot hope to find a larger one in general. Finding the exact base of the logarithm is one of the most difficult and important problems in Ramsey theory, which have recently witnessed a major breakthrough by the work of Campos, Griffiths, Morris, and Sahasrabudhe [Campos2023], which was subsequently optimized by Gupta, Ndiaye, Norin, and Wei [Gupta2024].
To obtain a bigger than logarithmically large homogeneous set, it is necessary to put additional assumptions on the given graphs. A class of graphs $\mathcal{C}$ is said to have the Erdős-Hajnal property if there is a constant $e>0$ such that each graph $G\in\mathcal{C}$ on $n$ vertices has a homogeneous subset of size $n^{e}$. The celebrated Erdős-Hajnal conjecture states that for each graph $H$ the class of $H$-free graphs admits the Erdős-Hajnal property [Erdos1977]. The conjecture is widely open in full, and was only recently resolved for all graphs on $5$ vertices [Nguyen2023][Chudnovsky2013].
A more restrictive property for a graph class than being $H$-free is having bounded VC-dimension. The Vapnik–Chervonenkis dimension, VC-dimension for short, is a measure of complexity of combinatorial structures, originally invented in the context of learning theory [Vapnik1971], which later proved to be fruitful in combinatorics as well. Structures with bounded VC-dimension often possess better Ramsey properties than general graphs [Fox2021][Janzer2024][Balogh2024]. In particular, Nguyen, Scott, and Seymour recently proved that graphs of bounded VC-dimensions admit the Erdős-Hajnal property [NSS_2023].
In this paper, we consider the following Ramsey statement by Ding, Oporowski, Oxley, and Vertigan [Ding_1996], independently proved using a similar method by Gravier, Maffray, Renault, and Trotignon [Gravier2004] building upon a result of Bauslaugh [Bauslaugh2001].
Theorem 1 ([Ding_1996][Gravier2004]).
There is a monotone unbounded function $f:\mathbb{N}\to\mathbb{N}$ such that the following holds. Let $G=(U,V,E)$ be a bipartite graph such that the part $B$ has $n$ vertices and no twins. Then it contains an induced subgraph $F=(U^{\prime},V^{\prime},E^{\prime})$ with $|U^{\prime}|=|V^{\prime}|\geq f(n)$ of $G$ that is isomorhic to either:
@@
Here the term twins refers to a pair of vertices with the same set of neighbors. All the graphs from the conclusion are bipartite with parts $u_{1},\dots,u_{n}$ and $v_{1},\dots,v_{n}$. The matching is the graph with $u_{i}\sim v_{j}$ iff $i=j$. The co-matching is the bipartite complement of a matching, i.e. $u_{i}\sim v_{j}$ iff $i\not=j$. The half-graph is the graph with $u_{i}\sim v_{j}$ iff $i\leq j$. Note that the binary complement, i.e. $u_{i}\sim v_{j}$ iff $i>j$, of the half-graph of order $n$ contains a half-graph of order $n-1$ as an induced subgraph.
Ding et al. proved Theorem 1 with log-log dependency of the size of $H$ on the size of $G$, and Gravier et al. showed that we cannot hope for better than logarithmic dependency. Inspired by the recent progress on the Erdős-Hajnal property, Dreier and Toruńczyk asked, [^1] In an open problem session within Algomanet Workshop in September 2024, Warsaw. whether Theorem 1 holds with a polynomial bound under the additional assumption that the graph $G$ has bounded VC-dimension.
@@
For every $d\in\mathbb{N}$, there is $c>0$ such that the following holds. Let $G=(U,V,E)$ be a bipartite graph of VC-dimension at most $d$ such that the part $V$ has $n$ vertices and no twins. Then it contains an induced subgraph $F=(U^{\prime},V^{\prime},E^{\prime})$ with $|U^{\prime}|=|V^{\prime}|\in\Omega_{d}(n^{c})$ that is isomorhic to either:
@@
Moreover, since Theorem 1 plays a role in model theory, where it is often considered to be a folklore result [Bonnet2024][Gajarsky2023], we comment on further improvements of Theorem 2 within model-theoretic context in Section 4. In fact, the applications in model theory was the original motivation of this paper.
@@ 2 Preliminaries
All our graphs are simple and finite. All our matrices are binary. Following the paper of Ding et al. [Ding_1996], we mostly work in the language of biadjacency matrices. The biadjacency matrix of a bipartite graph $G=(U,V,E)$ is the matrix $A\in\{0,1\}^{|U|\times|V|}$ with $A_{i,j}=1$ iff $\{u_{i},v_{j}\}\in E$. Natural operations for biadjacency matrices include permutation and deletion of rows/columns. Note that a matrix $B$ that is obtained from $A$ by these operations is the biadjacency matrix of an induced subgraph of $G$. We call such a matrix $B$ a submatrix of $A$.
@@
Observe that the square $(\alpha,\beta,\gamma)$-matrices with $\alpha\not=\beta$ correspond to the biadjacency matrices of the graphs from the conclusion of Theorem 2 (and the binary complement of a half-graph). Consequently, our result can be stated in the matrix language as follows.
@@
For every $d\in\mathbb{N}$, there is $c>0$ such that the following holds. Let $A$ be a binary matrix of VC-dimension at most $d$ with at least $n$ columns, no two of which are identical. Then it contains a square $(\alpha,\beta,\gamma)$-submatrix with $\alpha\not=\beta$ of size $n^{c}\times n^{c}$.
The VC-dimension of a matrix is defined in the next section.
@@ 2.1 VC-dimension
The Vapnik–Chervonenkis dimension is a numeric parameter capturing the complexity of a set system [Vapnik1971].
@@
If $B$ is (the binary complement of) a submatrix of $A$ (or its transposition), then $\operatorname{VC-dim}(B)\leq\operatorname{VC-dim}(A)$.
@@ 2.2 Growth function
The classical Sauer-Shelah lemma states that systems of bounded VC-dimension have the growth function bounded by a polynomial [Sauer_1972][Shelah_1972].
Theorem 4 (Sauer-Shelah lemma [Sauer_1972][Shelah_1972]).
@@ 2.3 Erdős-Hajnal property
As stated in the introduction, we say that a class of graphs $\mathcal{C}$ has the Erdős-Hajnal property if the graphs from $\mathcal{C}$ admit polynomially large homogeneous sets. A recent theorem of Nguyen, Scott, and Seymour that verifies the Erdős-Hajnal property for graphs of bounded VC-dimension [NSS_2023] plays a crucial part in our proof.
Theorem 5 ([NSS_2023]).
For every $d\geq 1$, there exists $e(d)>0$ such that every graph $G$ on $n$ vertices of VC-dimension at most $d$ contains a homogeneous set of size at least $n^{e(d)}$.
@@ 3 Proof
Given a matrix of VC-dimension at most $d$ with at least $n$ columns, no two of which are identical, our goal is to find a square $(\alpha,\beta,\gamma)$-submatrix with $\alpha\not=\beta$ of polynomial size. The original proof of Ding et al. is carried out in two steps. First, they obtain a square $(\alpha,\beta,*)$-submatrix of the given matrix with $\alpha\not=\beta$, where * is a Joker (i.e. no restriction for entries below the diagonal). As the second step, they apply the Ramsey theorem to an auxiliary graph to homogenize the lower triangular matrix. Both steps shrink the matrix logarithmically.
In our case, we prove that the auxiliary graph constructed from a matrix of bounded VC-dimension has bounded VC-dimension as well (Lemma 2). This allows us to perform the second step from the $(\alpha,\beta,*)$-submatrix to the $(\alpha,\beta,\gamma)$-submatrix with polynomial dependency by Theorem 5.
@@ 3.1 Switch submatrix
The following important technical lemma says that in our setting of bounded VC-dimensions we have polynomially large switch submatrices. We present the lemma in greater generality for matrices whose column systems have the growth function bounded by a polynomial.
@@
Let $A$ be a matrix whose system of columns has the growth function $\pi(n)\leq bn^{a}$ with $a\geq 1,b>0$. Suppose $A$ has $n$ columns, no two of which are identical. Then $A$ contains a switch submatrix $B$ of size at least $zn^{1/a}$ for any $z<1/(2+b)$.
@@
Starting from the empty submatrix $B_{0}$, we greedily construct switch submatrices $B_{1},B_{2},\dots$, with $B_{i}$ having size $i$. Each $B_{i+1}$ is obtained by extending $B_{i}$ by a new row and two columns. We only need to prove that we can run the process for at least $zn^{1/a}$ steps.
@@
It remains to show that such a vector $x^{*}\in\{0,1\}^{\mathcal{R}}$ exists when extending the submatrix $B_{i}$ with $i<zn^{1/a}$. By the assumption on the growth function, we have at most $bi^{a}$ vectors $x\in\{0,1\}^{\mathcal{R}}$ for which the set $\Gamma(x)$ is non-empty. Therefore, the average size of a non-empty $\Gamma(x)\setminus\mathcal{C}$ is at least
$\frac{n-|\mathcal{C}|}{bi^{a}}=\frac{n-2i}{bi^{a}}>\frac{n-2zn^{1/a}}{b(zn^{1/a})^{a}}=\frac{1}{bz^{a}}\left(1-\frac{2z}{n^{1-1/a}}\right)\geq\frac{1}{bz}(1-2z)>1$
as we need. Note that the last inequality is equivalent with the assumption $z<1/(2+b)$. ∎
@@
Let $A$ be a matrix of VC-dimension at most $d$ with $n$ columns, no two of which are identical. Then $A$ contains a switch submatrix $B$ of size at least $\Omega_{d}(n^{1/d})$.
@@
By Theorem 4, the class of matrices of VC-dimension at most $d$ has the growth function bounded by $\Phi_{d}(n)\in\mathcal{O}_{d}(n^{d})$. The conclusion then follows from Lemma 1. ∎
@@ 3.2 Upper graph and its VC-dimension
Let $A\in\{0,1\}^{n\times n}$ be a square $(*,0,*)$-matrix and $G$ be the upper graph of $A$. For any $d\in\mathbb{N}$, we have that $\operatorname{VC-dim}(G)\geq 2d$ implies $\operatorname{VC-dim}(A)\geq d/2$.
Conversely, for any $d\in\mathbb{N}$, if $\operatorname{VC-dim}(A)\leq d$, then $\operatorname{VC-dim}(G)\leq 4d+1$.
@@
Let $B$ be the adjacency matrix of $G$. The VC-dimension of $G$ is at least $2d$, which is witnessed by a shattered set $X=\{v_{1}<\dots<v_{2d}\}\subseteq[n]$ of size $d$. Let $Y=\{u_{1}<\dots<u_{2^{2d}}\}\subseteq[n]$ be the set of vertices whose neighborhoods witness that $X$ is shattered.
We divide the set $X$ into $X_{1}=\{v_{1}<\dots<v_{d}\}$ and $X_{2}=\{v_{d+1}<\dots<v_{k}\}$. Moreover, we split the set $Y$ into $Y_{1}=\{u_{i}\in Y:u_{i}<v_{d}\}$ and $Y_{2}=\{u_{i}\in Y:u_{i}\geq v_{d}\}$. Certainly, one of the sets $Y_{1}$ or $Y_{2}$ has size at least $2^{2d-1}$. If $|Y_{2}|\geq 2^{2d-1}$, the submatrix $C$ of $B$ induced by the rows $X_{1}$ and columns $Y_{2}$ lies entirely in the upper triangle of $B$, so it is a submatrix of $A$ as well. In the other case, if $|Y_{1}|\geq 2^{2d-1}$, the submatrix $D$ of $B$ induced by $X_{2}\times Y_{1}$ lies in the bottom triangle of $B$. By symmetry of $B$, the transposition of $D$ appears in the upper triangle of $B$ and hence also in $A$. The rest of the proof is then the same with the submatrix $D^{\top}$ instead of $C$ and with the role of rows and columns reversed.
[…]
$2^{d-1}\leq\text{ \#distinct patterns on $X_{1}$ }\leq\Phi_{\operatorname{VC-dim}(A)}(d)=\sum_{i=0}^{\operatorname{VC-dim}(A)}\binom{d}{i},$
which implies that $\operatorname{VC-dim}(A)\geq d/2$.
The part “Conversely” is the contraposition with correct rounding that includes odd values of $\operatorname{VC-dim}(G)$d-1. ∎

### v2

# A Polynomial Ramsey Statement for Bounded VC-dimension

@@ 1 Introduction
Ramsey theory is an area of mathematics asserting that each sufficiently large object contains a still large well-structured subobject. The foundational theorem, proved by Frank Ramsey in 1930, states that every sufficiently large graph contains a still large homogeneous (induced) subgraph, i.e. a clique or an independent set [30]. It is well known by a theorem of Erdős and Szekeres [19] that there is always a logarithmically large homogeneous subgraph. However, Erdős showed by a probabilistic construction that, in general, we cannot hope for larger than logarithmic [17]. Finding the exact base of the logarithm is one of the most difficult and important problems in Ramsey theory, which have recently witnessed a major breakthrough by the work of Campos, Griffiths, Morris, and Sahasrabudhe [10], which was subsequently optimized by Gupta, Ndiaye, Norin, and Wei [23].
To obtain a bigger than logarithmically large homogeneous subgraph, it is necessary to put additional assumptions on the given graphs. A class of graphs $\mathcal{C}$ is said to have the Erdős-Hajnal property if there is a constant $e>0$ such that each graph $G\in\mathcal{C}$ on $n$ vertices has a homogeneous subgraph of size $n^{e}$. The celebrated Erdős-Hajnal conjecture states that for each graph $H$ the class of $H$-free graphs admits the Erdős-Hajnal property [18]. The conjecture is widely open in full, and was only recently resolved for all graphs on $5$ vertices [28][11].
A more restrictive property for a graph class than being $H$-free is having bounded VC-dimension. The Vapnik–Chervonenkis dimension, VC-dimension for short, is a measure of complexity of combinatorial structures, originally invented in the context of learning theory [36], which later proved to be fruitful in combinatorics as well. Structures with bounded VC-dimension often possess better Ramsey properties than general graphs [20][24][2]. In particular, Nguyen, Scott, and Seymour recently proved that graphs of bounded VC-dimensions admit the Erdős-Hajnal property [27].
In this paper, we consider the following Ramsey statement by Ding, Oporowski, Oxley, and Vertigan [12], which was independently proved by Alekseev [1] (in Russian), and by Gravier, Maffray, Renault, and Trotignon [22] building upon a result of Bauslaugh [3].
Theorem 1 ([12][1][22]).
There is a monotone unbounded function $f:\mathbb{N}\to\mathbb{N}$ such that the following holds. Let $G=(U,V,E)$ be a bipartite graph such that the part $V$ has $n$ vertices and no twins. Then it contains an induced subgraph $F=(U^{\prime},V^{\prime},E^{\prime})$ with $|U^{\prime}|=|V^{\prime}|\geq f(n)$ that is isomorphic to either:
@@
Here the term twins refers to a pair of vertices with the same set of neighbors. All the graphs from the conclusion are bipartite with parts $u_{1},\dots,u_{n}$ and $v_{1},\dots,v_{n}$. The matching is the graph with $u_{i}\sim v_{j}$ iff $i=j$. The co-matching is the bipartite complement of a matching, i.e. $u_{i}\sim v_{j}$ iff $i\not=j$. The half-graph is the graph with $u_{i}\sim v_{j}$ iff $i\leq j$. Note that the binary complement, i.e. $u_{i}\sim v_{j}$ iff $i>j$, of the half-graph of order $n$ contains a half-graph of order n1n-1 as an induced subgraph.
Remarkably, all [12][1][22] proved Theorem 1 with log-log dependence of the size of $F$ on the size of $G$, and Gravier et al. showed that we cannot hope for better than logarithmic dependency [22]. In view of the recent progress on the Erdős-Hajnal property, it is a natural question, [^1] The author learned about the question from Jan Dreier and Szymon Toruńczyk in an open problem session within Algomanet Workshop in September 2024, Warsaw. whether Theorem 1 holds with a polynomial bound under the additional assumption that the graph $G$ has bounded VC-dimension.
@@
For every $d\in\mathbb{N}$, there is $c>0$ such that the following holds. Let $G=(U,V,E)$ be a bipartite graph of VC-dimension at most $d$ such that the part $V$ has $n$ vertices and no twins. Then it contains an induced subgraph $F=(U^{\prime},V^{\prime},E^{\prime})$ with $|U^{\prime}|=|V^{\prime}|\in\Omega(n^{c})$ that is isomorphic to either:
@@
Moreover, since Theorem 1 plays a role in model theory, where it is often considered to be a folklore result [6][21], we comment on further improvements of Theorem 2 within model-theoretic context in Section 4. In fact, the applications in model theory were the original motivation of this paper.
@@ 2 Preliminaries
All our graphs are simple and finite. All our matrices are binary. Following the paper of Ding et al. [12], we mostly work in the language of biadjacency matrices. The biadjacency matrix of a bipartite graph $G=(U,V,E)$ is the matrix $A\in\{0,1\}^{|U|\times|V|}$ with $A_{i,j}=1$ iff $\{u_{i},v_{j}\}\in E$. Natural operations for biadjacency matrices include permutation and deletion of rows/columns. Note that a matrix $B$ that is obtained from $A$ by these operations is the biadjacency matrix of an induced subgraph of $G$. We call such a matrix $B$ a submatrix of $A$.
@@
If not all the parameters $\alpha,\beta,\gamma$ are equal, we say that the $(\alpha,\beta,\gamma)$-matrix is inhomogeneous. Observe that the inhomogeneous square $(\alpha,\beta,\gamma)$-matrices correspond to the biadjacency matrices of the graphs from the conclusion of Theorem 2 (and the binary complement of a half-graph). Consequently, our result can be stated in the matrix language as follows.
@@
For every $d\in\mathbb{N}$, there is $c>0$ such that the following holds. Let $A$ be a binary matrix of VC-dimension at most $d$ with at least $n$ columns, no two of which are identical. Then it contains an inhomogeneous square $(\alpha,\beta,\gamma)$-submatrix of size $n^{c}\times n^{c}$.
In particular, we ensure that $\alpha\not=\beta$. The VC-dimension of a matrix is defined in the next section.
@@ 2.1 VC-dimension
The Vapnik–Chervonenkis dimension is a numeric parameter capturing the complexity of a set system [36].
@@
If $B$ is (the binary complement of) a submatrix of $A$ (or its transpose), then $\operatorname{VC-dim}(B)\leq\operatorname{VC-dim}(A)$.
@@ 2.2 Growth function
The classical Sauer-Shelah lemma states that systems of bounded VC-dimension have the growth function bounded by a polynomial [31][32].
Theorem 4 (Sauer-Shelah lemma [31][32]).
@@ 2.3 Erdős-Hajnal property
As stated in the introduction, we say that a class of graphs $\mathcal{C}$ has the Erdős-Hajnal property if the graphs from $\mathcal{C}$ admit polynomially large homogeneous subgraph. A recent theorem of Nguyen, Scott, and Seymour that verifies the Erdős-Hajnal property for graphs of bounded VC-dimension [27] plays a crucial part in our proof.
Theorem 5 ([27]).
For every $d\geq 1$, there exists $e(d)>0$ such that every graph $G$ on $n$ vertices of VC-dimension at most $d$ contains a homogeneous subgraph of size at least $n^{e(d)}$.
@@ 3 Proof
Given a matrix of VC-dimension at most $d$ with at least $n$ columns, no two of which are identical, our goal is to find an inhomogeneous square $(\alpha,\beta,\gamma)$-submatrix with $\alpha\not=\beta$ of polynomial size. The original proof of Ding et al. is carried out in two steps. First, they obtain a square $(\alpha,\beta,*)$-submatrix of the given matrix with $\alpha\not=\beta$, where * is a Joker (i.e. no restriction for entries below the diagonal). As the second step, they apply the Ramsey theorem to an auxiliary graph to homogenize the lower triangular matrix. Both steps shrink the matrix logarithmically.
In our case, we prove that the auxiliary graph constructed from a matrix of bounded VC-dimension has bounded VC-dimension as well (Lemma 2). This allows us to perform the second step from the $(\alpha,\beta,*)$-submatrix to the $(\alpha,\beta,\gamma)$-submatrix with a polynomial dependence by Theorem 5.
@@ 3.1 Switch submatrix
The following important technical lemma says that in our setting of bounded VC-dimensions we have polynomially large switch submatrices. We present the lemma for a general bounding function $f:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$ that satisfies some reasonable assumptions. That is, we assume that $f$ is convex with $f(0)=0$ and $f(x)\geq 2x$ for all $x\in\mathbb{R}^{+}_{0}$. Note that this implies $f$ to be strictly increasing (by Jensen’s inequality). Thus, $f$ has a well-defined (concave) inverse $f^{-1}$.
@@
Let $A$ be a matrix whose system of columns has a growth function $\pi\leq f$ with $f$ as above. Suppose $A$ has $n$ columns, no two of which are identical. Then $A$ contains a switch submatrix $B$ of size at least $\frac{1}{2}f^{-1}(n)$.
@@
Starting from the empty submatrix $B_{0}$, we greedily construct switch submatrices $B_{1},B_{2},\dots$, with $B_{i}$ having size $i$. Each $B_{i+1}$ is obtained by extending $B_{i}$ by a new row and two columns. We only need to prove that we can run the process for at least $\frac{1}{2}f^{-1}(n)$ steps.
@@
It remains to show that such a vector $x^{*}\in\{0,1\}^{\mathcal{R}}$ exists when extending the submatrix $B_{i}$ with $i<\frac{1}{2}f^{-1}(n)$. We have at most $\pi(i)$ vectors $x\in\{0,1\}^{\mathcal{R}}$ for which the set $\Gamma(x)$ is non-empty. Therefore, the average size of a non-empty $\Gamma(x)\setminus\mathcal{C}$ is at least
$\frac{n-|\mathcal{C}|}{\pi(i)}\geq\frac{n-2i}{f(i)}=\frac{n}{f(i)}-\frac{2i}{f(i)},$
using $\pi(i)\leq f(i)$ and $|\mathcal{C}|=2i$. Clearly, $\frac{2i}{f(i)}\leq 1$ due to the assumption $f(x)\geq 2x$. We proceed to lower-bound the other fraction. We have
$\frac{n}{f(i)}>\frac{n}{f(\frac{1}{2}f^{-1}(n))}=2\cdot\frac{\frac{1}{2}n}{f(\frac{1}{2}f^{-1}(n))},$
where the inequality follows from $i<\frac{1}{2}f^{-1}(n)$ and the fact that $f$ is strictly increasing. We claim that the last fraction is at least $1$. Indeed, this statement is equivalent to
$f\left(\frac{1}{2}f^{-1}(n)\right)\leq\frac{1}{2}n\qquad\Leftrightarrow\qquad\frac{1}{2}f^{-1}(n)\leq f^{-1}\left(\frac{1}{2}n\right)\qquad\Leftrightarrow\qquad\frac{f^{-1}(0)+f^{-1}(n)}{2}\leq f^{-1}\left(\frac{0+n}{2}\right),$
using that $f^{-1}(0)=0$, where the last inequality is exactly the Jensen inequality applied to the concave function $f^{-1}$. To conclude, we obtain
$\frac{n-|\mathcal{C}|}{\pi(i)}\geq\frac{n}{f(i)}-\frac{2i}{f(i)}>2\cdot 1-1=1,$
and therefore, $|\Gamma(x^{*})\setminus\mathcal{C}|\geq 2$. ∎
@@
Let $A$ be a matrix of VC-dimension at most $d$ with $n$ columns, no two of which are identical. Then $A$ contains a switch submatrix $B$ of size $\Omega(n^{1/d})$.
@@
By Theorem 4, the class of matrices of VC-dimension at most $d$ has the growth function bounded by $\Phi_{d}(n)\in\mathcal{O}(n^{d})$. Clearly, a function $cn^{d}\geq\Phi_{d}(n)$ with $c\geq 2,d\geq 1$ satisfies all the assumptions on the function $f$ from Lemma 1. Hence, we obtain a switch submatrix of $B$ of the size $\frac{1}{2c}n^{1/d}$ by Lemma 1. ∎
@@ 3.2 Upper graph and its VC-dimension
Let $A\in\{0,1\}^{n\times n}$ be a square (,0,)(*,0,*)-matrix and $G$ be the upper graph of $A$. For any $d\in\mathbb{N}$, we have that $\operatorname{VC-dim}(G)\geq 4d$ implies $\operatorname{VC-dim}(A)\geq d$.
Therefore, if $\operatorname{VC-dim}(A)\leq d$, then $\operatorname{VC-dim}(G)\leq 4d+3$.
@@
Let $B$ be the adjacency matrix of $G$. The VC-dimension of $G$ is at least $4d$, which is witnessed by a shattered set $X=\{v_{1}<\dots<v_{4d}\}\subseteq[n]$ of size $4d$. Let $Y=\{u_{1}<\dots<u_{2^{4d}}\}\subseteq[n]$ be the set of vertices whose neighborhoods witness that $X$ is shattered.
We divide the set $X$ into $X_{1}=\{v_{1}<\dots<v_{2d}\}$ and $X_{2}=\{v_{2d+1}<\dots<v_{4d}\}$. Moreover, we split the set $Y$ into $Y_{1}=\{u_{i}\in Y:u_{i}<v_{2d}\}$ and $Y_{2}=\{u_{i}\in Y:u_{i}\geq v_{2d}\}$. Certainly, one of the sets $Y_{1}$ or $Y_{2}$ has size at least $2^{4d-1}$. If $|Y_{2}|\geq 2^{4d-1}$, the submatrix $C$ of $B$ induced by the rows $X_{1}$ and columns $Y_{2}$ lies entirely in the upper triangle of $B$, so it is a submatrix of $A$ as well. In the other case, if $|Y_{1}|\geq 2^{4d-1}$, the submatrix $D$ of $B$ induced by $X_{2}\times Y_{1}$ lies in the bottom triangle of $B$. By the symmetry of $B$, the transpose of $D$ appears in the upper triangle of $B$ and hence also in $A$. The rest of the proof is then the same with the submatrix $D^{\top}$ instead of $C$ and with the role of rows and columns reversed.
Suppose that $|Y_{2}|\geq 2^{4d-1}$. As we argued in the previous paragraph, the submatrix $C$ of $B$ induced by the rows $X_{1}$ and columns $Y_{2}$ is a submatrix of $A$ as well. We claim that the rows $X_{1}$ of $A$, in particular the submatrix $C$, witness that $\operatorname{VC-dim}(A)\geq d$. The matrix $C$ contains at least $2^{4d-1}/\,2^{2d}=2^{2d-1}$ distinct columns since it has at least $2^{4d-1}$ columns in total and any pattern may occur in at most $2^{|X_{2}|}=2^{2d}$ of them. From Theorem 4, we obtain the inequality
$2^{2d-1}\leq\text{ \#distinct patterns on $X_{1}$ }\leq\Phi_{\operatorname{VC-dim}(A)}(2d)=\sum_{i=0}^{\operatorname{VC-dim}(A)}\binom{2d}{i}$.
This implies $\operatorname{VC-dim}(A)\geq d$ as we have the equality $\Phi_{d}(2d)=2^{2d-1}$ by basic properties of binomial coefficients.
The part “Therefore” is the contrapositive that includes the values of $\operatorname{VC-dim}(G)$ that are not divisible by $4$. ∎